OpenStudy (anonymous):

What is the slope of the line passing through the points (2, 5) and (0, −4)

OpenStudy (mayankdevnani):

use previous formula !!!

OpenStudy (mayankdevnani):

can you show your work ?? @Archer3545

OpenStudy (anonymous):

2+5 over 0 - (-4)?

OpenStudy (linn99123):

Nope! Remember its y2 - y1/x2 - x1! Where: (2, 5) = x1, y1 (0,-4) = x2, y2! Plug it in and solve :D

OpenStudy (anonymous):

so is it 2x1, 5x1 and 0x2, 4x2?

OpenStudy (linn99123):

So it would be: 5 - (-4)/0 - 2 9/-2 -4.5

OpenStudy (linn99123):

No :) x1 and y1 represents the first coordinates... x2 and y2 represents the second set of coordinates.

OpenStudy (gorv):

(x1,y1)=(2,5) (x2,y2)=(0,-4)

OpenStudy (gorv):

slope=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

ok i think i get it

OpenStudy (anonymous):

slope = (y2 - y1) / (x2 - x1) (2,5)...x1 = 2 and y1 = 5 (0,-4)...x2 = 0 and y2 = -4 now sub slope = (-4 - 5) / (0 - 2) slope = -9/-2 slope = 9/2